OpenStudy (anonymous):

hi Solve the word problem. A group of roommates decide to split the cost of a stereo system for their house. If a is the amount spent on the stereo system and n is the number of roommates, the formula C = determines each person’s share of the cost. Each person’s share was $107.50 and 6 roommates live in the house. What was the cost of the stereo system? A. $113.50 B. $607.50 C. $642.00 D. $645.00

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

whoever helps me i will be there fan.

OpenStudy (anonymous):

4. A group of roommates decide to split the cost of a stereo system for their house. If a is the amount spent on the stereo system and n is the number of roommates, the formula C = a/n determines each person’s share of the cost. Each person’s share was $107.50 and 6 roommates live in the house. What was the cost of the stereo system? C = a/n Cn = a a = Cn = $107.50 * 6 = D. $645.00

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

medal and fan please